The driver of a car traveling on the highway suddenly slams on the brakes because of a slowdown in traffic ahead.

A) If the car’s speed decreases at a constant rate from 74 mi/h to 50 mi/h in 3.0 s, what is the magnitude of its acceleration, assuming that it continues to move in a straight line?

Answer is in mi/h^2

B) What distance does the car travel during the braking period?

Answer is in ft
Answer #1

here,

initial speed , u = 74 mi/h

final speed , v = 50 mi/h

time taken , t = 3 s = 1/1200 hr

let the accelration be a

using first equation of motion

v = u + a * t

50 = 74 + a * 1/1200

a = - 28800 mi/h^2

B)

let the distance travelled be s

using third equation of motion

v^2 - u^2 = 2 * a * s

50^2 - 74^2 = - 2 * 28800 * s

solving for s

s = 0.052 mile = 272.8 ft